# zkLoRA: Fine-Tuning Large Language Models with Verifiable Security via Zero-Knowledge Proofs

### v1

@@ Abstract
Fine-tuning large language models (LLMs) is crucial for adapting them to specific tasks, yet it remains computationally demanding and raises concerns about correctness and privacy, particularly in untrusted environments. Although parameter-efficient methods like Low-Rank Adaptation (LoRA) significantly reduce resource requirements, ensuring the security and verifiability of fine-tuning under zero-knowledge constraints remains an unresolved challenge. To address this, we introduce zkLoRA, the first framework to integrate LoRA fine-tuning with zero-knowledge proofs (ZKPs), achieving provable security and correctness. zkLoRA employs advanced cryptographic techniques—such as lookup arguments, sumcheck protocols, and polynomial commitments—to verify both arithmetic and non-arithmetic operations in Transformer-based architectures. The framework provides end-to-end verifiability for forward propagation, backward propagation, and parameter updates during LoRA fine-tuning, while safeguarding the privacy of model parameters and training data. Leveraging GPU-based implementations, zkLoRA demonstrates practicality and efficiency through experimental validation on open-source LLMs like LLaMA, scaling up to 13 billion parameters. By combining parameter-efficient fine-tuning with ZKPs, zkLoRA bridges a critical gap, enabling secure and trustworthy deployment of LLMs in sensitive or untrusted environments.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2026.242361 www.ndss-symposium.org
@@ I Introduction
The rapid advancement of large language models (LLMs) has transformed natural language processing (NLP), enabling breakthroughs in tasks such as text generation, summarization, and machine translation. However, fine-tuning these massive models for specific tasks remains a significant challenge due to their enormous parameter sizes and associated computational costs [1, 2]. To address this, techniques like Low-Rank Adaptation (LoRA) [3] have emerged, offering a parameter-efficient way to fine-tune LLMs by introducing trainable low-rank matrices that adapt pretrained weights without modifying the base model. LoRA achieves high performance with significantly reduced resource requirements, making it a practical solution for many applications.
Despite these advancements, the security and correctness of fine-tuning processes remain an open problem, especially in scenarios involving sensitive data or untrusted environments [4, 5]. For example, when fine-tuning is conducted on proprietary models or outsourced to third-party platforms, there is a critical need to ensure that the computations are performed correctly without exposing the model parameters or training data. Existing research has explored zero-knowledge proofs (ZKPs) to verify computations in cryptographic and machine learning contexts, such as inference or simple training tasks. Frameworks like zkML [6] and zkLLM [7] have demonstrated the feasibility of applying ZKPs to machine learning, but these works primarily focus on inference or small-scale models. The problem of achieving zero-knowledge verifiability for fine-tuning large-scale LLMs, particularly under parameter-efficient methods like LoRA, remains unexplored and presents significant technical challenges.
@@
To the best of our knowledge, there is no existing work that provides zero-knowledge verifiability for the fine-tuning of large-scale LLMs using LoRA or similar parameter-efficient methods. Addressing this gap is both technically challenging and crucial for enabling secure and trustworthy deployment of LLMs in sensitive or trustless environments. To tackle these challenges, we propose zkLoRA, a novel framework that integrates LoRA fine-tuning with zero-knowledge proofs to achieve verifiable security and correctness during the fine-tuning of LLMs. The key contributions and novelties of this work are summarized as follows:
@@
    First Zero-Knowledge Verifiable Fine-Tuning Framework: zkLoRA is the first framework to provide end-to-end zero-knowledge verifiability for the fine-tuning of large-scale LLMs. It ensures that all computational steps in LoRA fine-tuning, including forward propagation, backward propagation, and parameter updates, are provably correct without revealing sensitive information such as model parameters or training data.
@@
    Innovative Handling of Non-Arithmetic Operations: zkLoRA addresses the challenge of verifying non-arithmetic operations occurring in the forward and backward computation process through Transformer layers of LLMs. By leveraging lookup-based arguments, zkLoRA encodes these operations into zero-knowledge proof systems, making them verifiable under stringent privacy constraints.
@@
    Experimental Validation on Large-Scale Models: We demonstrate the practicality of zkLoRA by applying it to fine-tune open-source LLMs, such as LLaMA-2, using real-world datasets. The experimental results showcase the framework’s ability to achieve verifiable fine-tuning with resaonable computational overhead, paving the way for more secure and trustworthy applications of LLMs.
@@ II Preliminary
This section introduces the foundational background theories and techniques employed by zkLoRA.
@@ II-A1 Proofs and Arguments
For an NP language $L\subseteq\{0,1\}^{\ast}$ with an associated relation $R=\{(x,w)\}$, an interactive proof system is a pair of probabilistic polynomial-time (PPT) algorithms $(\mathcal{P},\mathcal{V})$ that satisfy the following requirements. Completeness: If the statement $x$ is true, we have $\Pr\left[\langle\mathcal{P}(x,w),\mathcal{V}(x)\rangle=1\right]\geq 1-\mathsf{negl}(\lambda)$ for every honest prover that follows the protocol properly. Soundness: If the statement $x$ is false, we have $\Pr\left[\langle\mathcal{P}(x,w),\mathcal{V}(x)\rangle=1\right]\leq 1-\mathsf{negl}(\lambda)$ for any PPT prover that can deviate from the protocol.
We say that an interactive proof is public coin if $\mathcal{V}$’s challenge in each round is independent of $\mathcal{P}$’s messages in previous rounds. Originally introduced by [8], these systems allow resource-constrained verifiers to efficiently validate claims, such as NP-hard computations, while maintaining privacy through mechanisms like zero-knowledge proofs.
Succinct Non-Interactive Arguments of Knowledge (SNARKs) are widely utilized in argument systems [9, 10, 11, 12, 13, 8]. They facilitate efficient verification of complex computations, such as those involving large circuits or datasets, by providing sublinear proof sizes and constant-time verification.
@@ II-A2 Polynomial Commitments
A polynomial commitment scheme enables a prover to commit to a polynomial $f$ of degree at most $d$ over a finite field $\mathbb{F}$, such that the prover can later open the evaluation $y=f(x)$ at any point $x\in\mathbb{F}$ with a succinct proof. Representative realizations include the Hyrax commitment [14] and its multi-opening extensions [15]. A commitment scheme must satisfy two essential security properties:
@@ II-B1 Sumcheck Protocol
The sumcheck protocol [16] is a fundamental component of modern interactive proof systems, facilitating verifiable computation of multivariate polynomial summations with sublinear complexity. Formally, given a polynomial $f\colon\mathbb{F}^{l}\rightarrow\mathbb{F}$ of degree $d$ in each variable, the protocol enables a verifier to confirm the correctness of a prover’s claimed sum: $S=\sum_{b_{1},\ldots,b_{l}\in\{0,1\}}f(b_{1},\ldots,b_{l})$ with communication and verification costs scaling logarithmically with size of $2^{l}$.
The sumcheck protocol proceeds through $l$ iterative rounds: in each round, the prover sends a univariate polynomial $g_{i}(x_{i})$, purportedly representing the partial sum of $f$ over all remaining variables, while the verifier checks consistency by sampling a random challenge $r_{i}\in\mathbb{F}$. After $l$ rounds, the sumcheck protocol reduces the verification task to evaluating $f$ at a single random point. The Schwartz-Zippel lemma [17, 18] underpins this probabilistic check, confining the soundness error to $\frac{d}{|\mathbb{F}|}$, where $d$ is the degree of the polynomial and $|\mathbb{F}|$ is the cardinality of the finite field.
However, the original sumcheck protocol is not zero-knowledge because $\mathcal{P}$ reveals evaluations of $f$. References [19, 20, 21] introduced zero-knowledge polynomial commitments to conceal $f$ in the interactive proof procedure of sumcheck, enabling zero-knowledge variants. Additionally, the Fiat-Shamir transform can convert the protocol into a non-interactive zero-knowledge proof under the random oracle model.
@@ II-B2 Multilinear Extensions
In practice, MLEs are indispensable for modern cryptographic protocols. For instance, in the sumcheck protocol, MLEs transform discrete summations over $\{0,1\}^{n}$ into polynomial evaluations over $\mathbb{F}^{n}$, enabling the verifier to efficiently validate complex claims via interactive proofs. Similarly, zk-SNARK frameworks such as Spartan [22] leverage MLEs to compress circuit constraints into multilinear forms, significantly reducing both proof size and verification overhead.
@@ II-C Verifiable Tensor Operation
Tensor operations (i.e., the computations using matrices) serve as foundational components of LLMs [23, 24, 1, 2]. When implemented in zero-knowledge proof systems, tensor operations fundamentally rely on the sumcheck protocol [16] and MLEs [25]. Consider dimensional-two tensors (i.e., matrices) $A\in\mathbb{F}^{D_{0}\times D_{1}}$, $B\in\mathbb{F}^{D_{1}\times D_{2}}$, and $C\in\mathbb{F}^{D_{0}\times D_{2}}$ that are defined over a finite field $\mathbb{F}$, where $C=A\cdot B$ represents the standard matrix product:
@@ II-D Lookup Arguments for Non-arithmetic Operations
In our work, we extend the lookup arguments [7], as commonly employed to achieve non-arithmetic operations within the domain of zero-knowledge proofs [26], specially incorporate it into zero-knowledge verifiable deep learning inference [27] and LLM inference [7], to the domain of gradient computation in LLMs. Lookup arguments are cryptographic primitives designed to efficiently prove set-membership conditions, enabling the prover to demonstrate to a verifier the correctness of non-arithmetic operation in LLMs. Formally, the prover holds a secret table $S=\{s_{1},s_{2},\ldots,s_{D}\}\subseteq\mathbb{F}^{D}$, where $\mathbb{F}$ denotes a finite field. The lookup arguments enable a prover to prove to a verifier that each element $s_{i}$ in $S$ is contained within a publicly known, predefined lookup table $T=\{t_{1},t_{2},\ldots,t_{N}\}\subseteq\mathbb{F}^{N}$.
@@
where each $m_{j}$ indicates the multiplicity of the corresponding table element $t_{j}$ within the secret tensor $S$. According to zkLLM [7], the membership condition $S\subseteq T$ can be equivalently represented in inner product form:
@@
A complete description of the procedure to prove ${S}\subset{T}$ is found in Protocol 1. In particular, in Line 1, $\mathsf{Setup}(T)$ generates a short witness $\llbracket T\rrbracket$ to a prescribed table $T$ known to both of the prover and the verifier; in Line 4, the prover constructs m{m} based on a tensor S{S} and table T{T} and commits to S{S} and m{m} using $\mathsf{Prep}({S},{T})$; finally, in Line 9, $\langle\mathcal{P},\mathcal{V}\rangle.\mathsf{Prove}(\llbracket{S}\rrbracket,\llbracket{m}\rrbracket,\llbracket T\rrbracket)$ is the interactive process between the prover and the verifier, proving that a secret tensor S{S} is element-wisely in $T$, which has been committed as $\llbracket T\rrbracket$.
@@
Require: The prover $\mathcal{P}$ knows $S\in\mathbb{F}^{D}$, and the prover $\mathcal{P}$ and the verifier $\mathcal{V}$ both know $S\in\mathbb{F}^{N}$, where $N,D$ are both powers of $2$ such that $N$ divides $D$.
@@
2:  return $\llbracket T\rrbracket\leftarrow\mathsf{Commit}(T;0)$ $\triangleright$ No hiding required
@@
13:  $\mathcal{P}$ and $\mathcal{V}$ run the sumcheck on (7), followed by the proofs of evaluation on $\llbracket A\rrbracket,\llbracket B\rrbracket,\llbracket S\rrbracket,\llbracket m\rrbracket,\llbracket T\rrbracket$
@@ III-A Structure of LLM
A LLM, such as LLaMA [2], is a specialized deep neural network architecture designed to map input token sequences to contextualized representations through a stack of Transformer layers. Fig. 1 illustrates a typical LLM architecture, which consists of $L$ cascaded Transformer layers followed by a final linear projection layer. In this subsection, we will introduce the specific structures and computations involved in the Transformer layers and the final output layer separately.

### v3

# VeriLoRA: Fine-Tuning Large Language Models with Verifiable Security via Zero-Knowledge Proofs

@@ Abstract
Fine-tuning large language models (LLMs) is crucial for adapting them to specific tasks, yet it remains computationally demanding and raises concerns about correctness and privacy, particularly in untrusted environments. Although parameter-efficient methods like Low-Rank Adaptation (LoRA) significantly reduce resource requirements, ensuring the security and verifiability of fine-tuning under zero-knowledge constraints remains an unresolved challenge. To address this, we introduce VeriLoRA, the first framework to integrate LoRA fine-tuning with zero-knowledge proofs (ZKPs), achieving provable security and correctness. VeriLoRA employs advanced cryptographic techniques—such as lookup arguments, sumcheck protocols, and polynomial commitments—to verify both arithmetic and non-arithmetic operations in Transformer-based architectures. The framework provides end-to-end verifiability for forward propagation, backward propagation, and parameter updates during LoRA fine-tuning, while safeguarding the privacy of model parameters and training data. Leveraging GPU-based implementations, VeriLoRA demonstrates practicality and efficiency through experimental validation on open-source LLMs like LLaMA, scaling up to 13 billion parameters. By combining parameter-efficient fine-tuning with ZKPs, VeriLoRA bridges a critical gap, enabling secure and trustworthy deployment of LLMs in sensitive or untrusted environments.
@@ I Introduction
The rapid advancement of large language models (LLMs) has transformed natural language processing (NLP), enabling breakthroughs in tasks such as text generation, summarization, and machine translation [ray2023chatgpt, zhang2022optopenpretrainedtransformer, touvron2023llama, 11164571, 11117175]. However, fine-tuning these massive models for specific tasks remains a significant challenge due to their enormous parameter sizes and associated computational costs [chowdhery2023palm, touvron2023llama]. To address this, techniques like Low-Rank Adaptation (LoRA) [hu2022lora] have emerged, offering a parameter-efficient way to fine-tune LLMs by introducing trainable low-rank matrices that adapt pretrained weights without modifying the base model. LoRA achieves high performance with significantly reduced resource requirements, making it a practical solution for many applications.
Despite these advancements, the security and correctness of fine-tuning processes remain an open problem, especially in scenarios involving sensitive data or untrusted environments [ghodsi2017safetynets, zhao2021veriml]. For example, when fine-tuning is conducted on proprietary models or outsourced to third-party platforms, there is a critical need to ensure that the computations are performed correctly without exposing the model parameters or training data [11082313]. Existing research has explored zero-knowledge proofs (ZKPs) to verify computations in cryptographic and machine learning contexts, such as inference or simple training tasks. Frameworks like zkML [chen2024zkml] and zkLLM [sun2024zkllm] have demonstrated the feasibility of applying ZKPs to machine learning, but these works primarily focus on inference or small-scale models. The problem of achieving zero-knowledge verifiability for fine-tuning large-scale LLMs, particularly under parameter-efficient methods like LoRA, remains unexplored and presents significant technical challenges.
@@
To the best of our knowledge, there is no existing work that provides zero-knowledge verifiability for the fine-tuning of large-scale LLMs using LoRA or similar parameter-efficient methods. Addressing this gap is both technically challenging and crucial for enabling secure and trustworthy deployment of LLMs in sensitive or trustless environments. To tackle these challenges, we propose VeriLoRA, a novel framework that integrates LoRA fine-tuning with zero-knowledge proofs to achieve verifiable security and correctness during the fine-tuning of LLMs. The key contributions and novelties of this work are summarized as follows:
@@
    First Zero-Knowledge Verifiable Fine-Tuning Framework: VeriLoRA is the first framework to provide end-to-end zero-knowledge verifiability for the fine-tuning of large-scale LLMs. It ensures that all computational steps in LoRA fine-tuning, including forward propagation, backward propagation, and parameter updates, are provably correct without revealing sensitive information such as model parameters or training data.
@@
    Innovative Handling of Non-Arithmetic Operations: VeriLoRA addresses the challenge of verifying non-arithmetic operations occurring in the forward and backward computation process through Transformer layers of LLMs. By leveraging lookup-based arguments, VeriLoRA encodes these operations into zero-knowledge proof systems, making them verifiable under stringent privacy constraints.
@@
    Experimental Validation on Large-Scale Models: We demonstrate the practicality of VeriLoRA by applying it to fine-tune open-source LLMs, such as LLaMA-2, using real-world datasets. The experimental results showcase the framework’s ability to achieve verifiable fine-tuning with resaonable computational overhead, paving the way for more secure and trustworthy applications of LLMs.
The structure of the remainder of this paper is as follows. Section II provides preliminary knowledge of zero knowledge. Section III outlines the typical structure of LLMs. Section IV details the design of VeriLoRA protocol. Section V analyzes the security of VeriLoRA protocol. Section VI presents our experimental setup and results. Section LABEL:related_work reviews related work. Section LABEL:conclusion concludes this work and discusses directions for future research.
@@ II Preliminary
This section introduces the foundational background theories and techniques employed by VeriLoRA.
@@ II-A1 Proofs and Arguments
For an NP language $L\subseteq\{0,1\}^{\ast}$ with an associated relation $R=\{(x,w)\}$, an interactive proof system is a pair of probabilistic polynomial-time (PPT) algorithms $(\mathcal{P},\mathcal{V})$, which are termed the prover and verifier, respectively. The protocol of $(\mathcal{P},\mathcal{V})$ must satisfy the following two requirements. Completeness: If the statement $x$ is true (meaning a valid witness $w$ exists and is known to an honest prover), we have $\Pr\left[\langle\mathcal{P}(x,w),\mathcal{V}(x)\rangle=1\right]\geq 1-\mathsf{negl}(\lambda)$ for every honest prover $\mathcal{P}$ that follows the protocol properly. Soundness: If the statement $x$ is false ($x\notin L$), we have $\Pr\left[\langle\mathcal{P}^{*}(x,w),\mathcal{V}(x)\rangle=1\right]\leq\mathsf{negl}(\lambda)$ for any PPT prover $\mathcal{P}^{*}$, even one that deviates maliciously from the protocol.
We say that an interactive proof is public coin if the verifier’s challenge in each round is independent of prover’s messages in previous rounds. Originally introduced by [goldwasser2019knowledge], these systems allow resource-constrained verifiers to efficiently validate claims, such as NP-hard computations, while maintaining privacy through mechanisms like zero-knowledge proofs.
Succinct Non-Interactive Arguments of Knowledge (SNARKs) are widely utilized in argument systems [gabizon2019plonk, chiesa2020marlin, gennaro2013quadratic, parno2016pinocchio, ben2014succinct, goldwasser2019knowledge]. They facilitate efficient verification of complex computations, such as those involving large circuits or datasets, by providing sublinear proof sizes and constant-time verification.
@@ II-A2 Polynomial Commitments
A polynomial commitment scheme enables a prover to commit to a polynomial $f$ of degree at most $d$ over a finite field $\mathbb{F}$, such that the prover can later open the evaluation $y=f(x)$ at any point $x\in\mathbb{F}$ with a succinct proof. Representative realizations include the Hyrax commitment [wahby2018doubly] and its multi-opening extensions [boneh2020efficient]. A commitment scheme must satisfy two essential security properties:
@@ II-B1 Sumcheck Protocol
The sumcheck protocol [cormode2012practical] is a fundamental component of modern interactive proof systems, facilitating verifiable computation of multivariate polynomial summations with sublinear complexity. Formally, given a polynomial $f\colon\mathbb{F}^{l}\rightarrow\mathbb{F}$ of degree $d$ in each variable, the protocol enables a verifier to confirm the correctness of a prover’s claimed sum: $S=\sum_{b_{1},\ldots,b_{l}\in\{0,1\}}f(b_{1},\ldots,b_{l})$ with communication and verification costs scaling logarithmically with size of $2^{l}$.
The sumcheck protocol proceeds through $l$ iterative rounds: in each round, the prover sends a univariate polynomial $g_{i}(x_{i})$, purportedly representing the partial sum of $f$ over all remaining variables, while the verifier checks consistency by sampling a random challenge $r_{i}\in\mathbb{F}$. After $l$ rounds, the sumcheck protocol reduces the verification task to evaluating $f$ at a single random point. The Schwartz-Zippel lemma [schwartz1980fast, zippel1979probabilistic] underpins this probabilistic check, confining the soundness error to $\frac{d}{|\mathbb{F}|}$, where $d$ is the degree of the polynomial and $|\mathbb{F}|$ is the cardinality of the finite field.
However, the original sumcheck protocol is not zero-knowledge because the prover reveals evaluations of $f$. References [xie2019libra, xie2022orion, chiesa2017zero] introduced zero-knowledge polynomial commitments to conceal $f$ in the interactive proof procedure of sumcheck, enabling zero-knowledge variants. Additionally, the Fiat-Shamir transform can convert the protocol into a non-interactive zero-knowledge proof under the random oracle model.
@@ II-B2 Multilinear Extensions
In practice, MLEs are indispensable for modern cryptographic protocols. For instance, in the sumcheck protocol, MLEs transform discrete summations over $\{0,1\}^{n}$ into polynomial evaluations over $\mathbb{F}^{n}$, enabling the verifier to efficiently validate complex claims via interactive proofs. Similarly, zk-SNARK frameworks such as Spartan [setty2020spartan] leverage MLEs to compress circuit constraints into multilinear forms, significantly reducing both proof size and verification overhead.
@@ II-C Verifiable Tensor Operation
Tensor operations (i.e., the computations using matrices) serve as foundational components of LLMs [radford2019language, achiam2023gpt, chowdhery2023palm, touvron2023llama]. When implemented in zero-knowledge proof systems, tensor operations fundamentally rely on the sumcheck protocol [cormode2012practical] and MLEs [lund1992algebraic]. Consider dimensional-two tensors (i.e., matrices) $A\in\mathbb{F}^{D_{0}\times D_{1}}$, $B\in\mathbb{F}^{D_{1}\times D_{2}}$, and $C\in\mathbb{F}^{D_{0}\times D_{2}}$ that are defined over a finite field $\mathbb{F}$, where $C=A\cdot B$ represents the standard matrix product:
@@ II-D Lookup Arguments for Non-arithmetic Operations
In our work, we extend the lookup arguments [sun2024zkllm], as commonly employed to achieve non-arithmetic operations within the domain of zero-knowledge proofs [thaler2022proofs], specially incorporate it into zero-knowledge verifiable deep learning inference [kang2022scaling] and LLM inference [sun2024zkllm], to the domain of gradient computation in LLMs. Lookup arguments are cryptographic primitives designed to efficiently prove set-membership conditions, enabling the prover to demonstrate to a verifier the correctness of non-arithmetic operation in LLMs. Formally, the prover holds a secret table $S=\{s_{1},s_{2},\ldots,s_{D}\}\subseteq\mathbb{F}^{D}$, where $\mathbb{F}$ denotes a finite field. The lookup arguments enable a prover to prove to a verifier that each element $s_{i}$ in $S$ is contained within a publicly known, predefined lookup table $T=\{t_{1},t_{2},\ldots,t_{N}\}\subseteq\mathbb{F}^{N}$.
@@
where each $m_{j}$ indicates the multiplicity of the corresponding public table element $t_{j}$ within the secret table $S$: $m_{j}=|\{i:s_{i}=t_{j}\}|$ for $i\in\{1,2,\cdots,N\}$. According to zkLLM [sun2024zkllm], the membership condition $S\subseteq T$ can be equivalently represented in inner product form:
@@
A complete description of the procedure to prove ${S}\subseteq{T}$ is found in Protocol 1. In particular, in Line 1, $\mathsf{Setup}(T)$ generates a short witness $\llbracket T\rrbracket$ to a prescribed table $T$ known to both of the prover and the verifier; in Line 4, the prover constructs m{m} based on S{S} and T{T} and commits to S{S} and m{m} using $\mathsf{Prep}({S},{T})$; finally, in Line 9, $\langle\mathcal{P},\mathcal{V}\rangle.\mathsf{Prove}(\llbracket{S}\rrbracket,\llbracket{m}\rrbracket,\llbracket T\rrbracket)$ is the interactive process between the prover and the verifier, proving that a secret tensor S{S} is element-wisely in $T$, which has been committed as $\llbracket T\rrbracket$.
@@
Require: The prover $\mathcal{P}$ knows $S\in\mathbb{F}^{D}$, and the prover $\mathcal{P}$ and the verifier $\mathcal{V}$ both know $T\in\mathbb{F}^{N}$, where the integers $N,D$ are both powers of $2$ such that $N$ divides $D$.
@@
2:  return $\llbracket T\rrbracket\leftarrow\mathsf{Commit}(T)$
@@
13:  $\mathcal{P}$ and $\mathcal{V}$ run sumcheck on (7), followed by the proofs of evaluation on $\llbracket A\rrbracket,\llbracket B\rrbracket,\llbracket S\rrbracket,\llbracket m\rrbracket,\llbracket T\rrbracket$
@@ III-A Structure of LLM
A LLM, such as LLaMA [touvron2023llama], is a specialized deep neural network architecture designed to map input token sequences to contextualized representations through a stack of Transformer layers. Fig. 1 illustrates a typical LLM architecture, which consists of $L$ cascaded Transformer layers followed by a final linear projection layer. In this subsection, we will introduce the specific structures and computations involved in the Transformer layers and the final output layer separately.